### v1

# The Flavor Composition of Supernova Neutrinos

Abstract
We show that standard matter effects in the outer layers of core-collapse supernovae significantly constrain the flavor composition of the neutrino flux, even with the enormous uncertainties originating from self-induced flavor conversions in the supernova core. Under certain conditions, the neutrino flux resulting from self-induced conversions can be represented as a combination of flavor eigenstates in an arbitrary flavor ratio configuration. In this scenario, we find that, for the normal mass ordering, the fraction of neutrinos with electron flavor reaching the Earth, denoted as $f_{\nu_{e}}$, is constrained to be less than $0.5$ for all energies throughout the emission phase, whereas, for inverted mass ordering, we anticipate neutrinos arriving in near flavor equipartition ($f_{\nu_{e}}\approx 1/3$). In case adiabaticity is violated in the region of standard matter effects, the result is flavor equipartition for both mass orderings. Subsequently, we elaborate on the impact of wave-packet decoherence during self-induced conversions and explore alternative scenarios that could affect the aforementioned results.
Introduction.— During a core-collapse supernova (SN) event, the immense gravitational binding energy of the progenitor star is predominantly liberated through a rapid emission of neutrinos, typically lasting around 10 seconds Colgate:1966ax ; Arnett:1966 . After being generated in the SN core, these neutrinos traverse the mantle and envelope, where they are thought to be pivotal in driving the explosion Wilson:1985 ; Bethe:1985sox .
Figure 1: Allowed regions for the flavor composition of supernova neutrinos arriving at the Earth for the cases of normal mass ordering (left) and inverted mass ordering (right). The colors represent three different scenarios. Red: The flux emerging from collective effects is composed of flavor eigenstates in an arbitrary flavor ratio configuration, and propagation in the outer layers is adiabatic, $f_{\nu_{e}}^{NO}\lesssim 0.5$ and $f_{\nu_{e}}^{IO}\approx 1/3$. Maroon (stripes): Flux emerging from collective effects in an arbitrary flavor ratio configuration, but non-adiabatic propagation through the $H$-resonance (presence of shock wave), $f_{\nu_{e}}^{NO}\approx f_{\nu_{e}}^{IO}\approx 1/3$. Pink: Flux of matter eigenstates in arbitrary amounts emerging from collective effects (for instance, due to wave packet decoherence), $f_{\nu_{e}}^{NO}\approx f_{\nu_{e}}^{IO}\lesssim 0.7$. In this figure, we consider the $3\sigma$ range of variability only for the parameters $\theta_{12}$ and $\theta_{13}$, which affect the fraction of electron flavor $f_{\nu_{e}}$. Analogous results for antineutrinos are presented in the Appendix.
In 1987, the detection of a few tens of SN neutrinos coming from the Large Magellanic Cloud confirmed the basic features of the core-collapse models, such as the mean energy and the duration of the neutrino burst Kamiokande-II:1987idp ; Bionta:1987qt ; Alekseev:1988gp . Nevertheless, the intricacies of the SN physics remain a challenge, largely due to the uncertainties surrounding neutrino flavor transformations in the core Tamborra:2020cul ; Ehring:2023lcd ; Ehring:2023abs ; Nagakura:2023mhr . In such a neutrino-dense environment, the coherent forward scattering of neutrinos onto each other gives rise to collective effects that drive a nonlinear flavor evolution Pantaleone:1992eq ; Pantaleone:1992xh ; Duan:2010bg ; Mirizzi:2015eza ; Chakraborty:2016yeg . These self-induced conversions are referred to as “slow” or “fast” depending on the characteristic spatial scale in which they manifest. Slow self-induced conversions develop over a length scale governed by the vacuum oscillation frequency, which is on the order of a few kilometers for typical SN neutrino energies ($\sim 10$ MeV) Duan:2010bg ; Mirizzi:2015eza . These conversions take place in an intermediate zone situated $\sim 100$ km away from the neutrino decoupling region, but before the standard matter-induced resonant neutrino conversion occurs Wolfenstein:1977ue ; Mikheyev:1985zog ; Mikheev:1986wj ; Mikheev:1987jp . Conversely, fast self-induced conversions have characteristic scales as small as centimeters and could significantly influence the flavor evolution, energy spectrum, and angular distribution of each neutrino species near the decoupling region Sawyer:2005jk ; Sawyer:2008zs ; Sawyer:2015dsa ; Chakraborty:2016lct ; Izaguirre:2016gsx . The lack of a full solution to the quantum kinetic equations governing this behavior leads to unpredictable consequences for the SN dynamics Ehring:2023abs ; Nagakura:2023mhr and, ultimately, for the neutrino signal detected at the Earth Abbar:2024nhz .
Amidst this conundrum, it is natural to wonder if a meaningful improvement in our understanding of the core-collapse phenomenon and neutrino propagation will be possible, even after the observation of the next galactic SN. In this Letter, we show that many features of the neutrino flux coming to the Earth can be understood by adopting a completely agnostic approach to the flavor composition emerging from the SN core, where collective effects dominate. By considering all possible flavor ratio configurations before the onset of standard matter effects, we observe that matter effects alone impose significant constraints on the flavor composition of SN neutrinos. Assuming normal ordering (NO) of neutrino masses, we find that the fraction of electron neutrinos arriving at the Earth, denoted by $f_{\nu_{e}}^{NO}$, remains consistently below $0.5$ for all energies during the emission phase, and gradually converges to near flavor equipartition ($f_{\nu_{e}}^{NO}\approx 1/3$) [^1] In this Letter, “flavor equipartition” simply means $f_{\nu_{e}}=1/3$, since discrimination between $\nu_{\mu}$ and $\nu_{\tau}$ signals in current and next-generation detectors is challenging Akhmedov:2002zj . during the passage of shock waves. In the inverted ordering (IO) scenario, we predict flavor equipartition ($f_{\nu_{e}}^{IO}\approx 1/3$) as a generic outcome, which remains robust throughout the entire neutrino burst, even in the presence of shock waves. We specify the circumstances in which the flux resulting from collective effects is accurately described as consisting of flavor eigenstates. In addition, we explore the impact of wave-packet decoherence in self-induced conversions, along with other possible scenarios, on our results. If confirmed by the next SN observation, these results will not only validate our fundamental understanding of the SN dynamics and neutrino conversion in the outermost layers, but also elucidate crucial properties of the neutrino fluxes emerging from the region of self-induced conversions.
Standard Matter Effects.— Neutrino-neutrino interactions are expected to play a subdominant role in flavor evolution after the first few hundred kilometers from the neutrino decoupling region Tamborra:2020cul ; Mirizzi:2015eza . At the outermost layers, neutrino evolution is predominantly affected by the matter potential generated by the coherent forward scattering of neutrinos on electrons, protons and neutrons Wolfenstein:1977ue . In practice, however, only the charged current interactions between electrons and electron (anti)neutrinos influence neutrino oscillations, as neutral current interactions with any of the particles in the medium affect all neutrino flavors equally and can be factored out in the evolution. Consequently, the effective Hamiltonian in the outer SN regions can be written in the flavor basis as follows
$\mathcal{H}_{\nu}=\frac{1}{2E}\,U\left(\begin{array}[]{ccc}0&0&0\\ 0&\Delta m_{21}^{2}&0\\ 0&0&\Delta m_{31}^{2}\end{array}\right)U^{\dagger}+\left(\begin{array}[]{ccc}V_{e}&0&0\\ 0&0&0\\ 0&0&0\end{array}\right)~{}.$ (1)
Here, $U$ represents the Pontecorvo–Maki–Nakagawa– Sakata (PMNS) matrix, $\Delta m^{2}_{ij}$ stands for the mass-squared differences and $E$ denotes the neutrino energy. The matter potential affects only the $\nu_{e}$ flavor and has the form $V_{e}=\sqrt{2}G_{F}\rho_{m}Y_{e}/m_{N}$, where $G_{F}$ is the Fermi constant, $Y_{e}$ is the electron number fraction, $m_{N}$ is the nucleon mass and $\rho_{m}$ is the local matter density. It is important to consider the implicit dependence of $V_{e}$ and $\mathcal{H}_{\nu}$ on time and space, as each neutrino follows a different trajectory inside the SN, and the physical conditions along each trajectory change with time. For antineutrinos, the matter potential has the opposite sign, $\bar{V}_{e}=-V_{e}$. Thus, the equation of motion for neutrinos in the SN outer layers is given by
$i\frac{d\rho_{\nu}}{dt}=[\mathcal{H}_{\nu},\rho_{\nu}]~{},$ (2)
where $\rho_{\nu}$ is the density matrix describing the neutrino flavor state. Observe that Eq. 2 is a simplified version of the complete expression that takes into account neutrino-neutrino interactions, advection, and collisions Sigl:1993ctk ; Cardall:2007zw .
As neutrinos propagate towards vacuum and the matter density ($\rho_{m}$) gradually decreases, they encounter two resonance layers where effective flavor conversions can occur Mikheyev:1985zog ; Mikheev:1986wj ; Mikheev:1987jp . These so-called $H$- and $L$-resonances occur when the following conditions are satisfied:
$\sqrt{2}G_{F}\frac{Y_{e}}{m_{N}}\rho_{m}^{\rm res}=\frac{\Delta m^{2}_{i1}\cos 2\theta_{1i}}{2E}~{},$ (3)
where ${\rm res}=H(L)$ corresponds to $i=3(2)$. For efficient flavor conversion, adiabaticity must be satisfied at the resonance layers. The degree of adiabaticity can be estimated by considering the parameter
$\gamma=\frac{\Delta m^{2}_{i1}\sin^{2}2\theta_{1i}}{2E\cos 2\theta_{1i}}\left(\frac{1}{Y_{e}\rho_{m}^{\rm res}}\left|\frac{d(Y_{e}\rho_{m})}{dr}\right|_{\rm res}\right)^{-1}~{},$ (4)
and propagation is adiabatic for $\gamma>1$ Dighe:1999bi .
[…]
Flavor eigenstates emerging from self-induced conversions.— Assuming an emission of $10^{58}$ neutrinos over the $10$-second duration of a galactic SN burst SajjadAthar:2021prg , at a distance of 10 kpc, we estimate the flux reaching the Earth to be approximately $10^{15}$ neutrinos per square meter per second. Before traveling to the Earth, all these neutrinos reach the boundary between collective effects and standard matter effects inside the SN with a quantum state described by a $3\times 3$ density matrix, where the diagonal elements in the flavor basis give the probability for the neutrino to be measured in one of the three flavors, while the off-diagonal elements indicate the ability of the neutrino to interfere with itself (quantum coherence) and generate flavor oscillations. The region of standard matter effects destroys any superposition of electron neutrinos with other flavors initially formed by collective effects, whereas the subspace $\nu_{\mu}-\nu_{\tau}$ retains its flavor superposition. As a result, the neutrino state either transitions to $\nu_{e}$ or to a superposition of $\nu_{\mu}$ and $\nu_{\tau}$.
The density matrix corresponding to the $\nu_{\mu}-\nu_{\tau}$ subspace (sometimes referred to as $\nu_{x}$) of the $k$-th neutrino exiting the collective effects region towards the Earth is given by
$\rho^{k}_{\nu_{x}}=\left(\begin{array}[]{cc}\alpha_{k}^{2}&\alpha_{k}\beta_{k}e^{-i\phi_{k}}\\ \alpha_{k}\beta_{k}e^{i\phi_{k}}&\beta_{k}^{2}\end{array}\right)~{}.$ (5)
Here, $\alpha_{k}$, $\beta_{k}$, and $\phi_{k}$ are real numbers, subject to the condition $\alpha_{k}^{2}+\beta_{k}^{2}=1$. Furthermore, $k$ ranges from $1$ to $N$, where $N$ represents the number of neutrinos with a specific energy leaving the collective effects region towards the Earth and having collapsed to non-electron flavors. The density matrix for the ensemble of $N$ neutrinos is
$\rho_{\nu_{x}}=\sum_{k=1}^{N}\left(\begin{array}[]{cc}\alpha_{k}^{2}&\alpha_{k}\beta_{k}e^{-i\phi_{k}}\\ \alpha_{k}\beta_{k}e^{i\phi_{k}}&\beta_{k}^{2}\end{array}\right)~{}.$ (6)
Since $N$ is very large, as mentioned at the beginning of this section, it is natural to assume that the off-diagonal elements of Eq. 6 average to zero, due to the many oscillatory terms composing the sum [^3] For simplicity, the density matrix in Eq. 5 corresponds to a pure state, namely |$\ket{\nu_{x}}=\alpha_{k}\ket{\nu_{\mu}}+\beta_{k}e^{i\phi_{k}}\ket{\nu_{\tau}}$. However, the same argument applies to states with an intermediate level of coherence, provided that the terms in the off-diagonal sum in Eq. 6 retain some degree of oscillatory behavior Kersten:2015kio . Later, we discuss situations where the assumption of vanishing off-diagonal terms is not valid.. Note that $\alpha_{k}$, $\beta_{k}$, and $\phi_{k}$ depend on the evolution of the $k$-th neutrino, which can be enormously diverse before entering the region of standard matter effects Hansen:2018apu . In such a situation, the superposition between $\nu_{\mu}$ and $\nu_{\tau}$ is effectively destroyed. Therefore, after collective effects, the neutrino flux density matrix becomes diagonal in the flavor basis and can be described as composed of the various flavor eigenstates in arbitrary proportions.
In what follows, the discussion is centered on neutrinos. For the results related to antineutrinos, see the Appendix.
Results: flavor ratios at the Earth.— In the SN region where neutrino-neutrino potentials are already subdominant and $\rho_{m}\gg\rho_{m}^{H}$, the mixing between the $\nu_{e}$ and the non-electron flavors is suppressed. As previously elucidated, $\nu_{e}$ can be approximated as $\nu_{3}^{m}$ $(\nu_{2}^{m})$ for NO (IO). Meanwhile, inside the subspace $\nu_{\mu}-\nu_{\tau}$, the effective mixing is near-maximal and $\nu_{\mu}$ and $\nu_{\tau}$ are both approximately equal mixtures of the same matter eigenstates: $\nu_{1}^{m}$ and $\nu_{2}^{m}$ ($\nu_{1}^{m}$ and $\nu_{3}^{m}$) for NO (IO). Consequently, even if the physical output of the central region of self-induced conversions are flavor eigenstates in random relative proportions, the effective mixing in the $\nu_{\mu}-\nu_{\tau}$ subspace ensures that two out of the three matter eigenstates consistently occur in approximately equal amounts.
Assuming adiabaticity of neutrino propagation, the relative proportions between the vacuum eigenstates that reach the Earth are identical to those of matter eigenstates propagating in the SN outer layers ($\nu_{1}^{m}\rightarrow\nu_{1}$, $\nu_{2}^{m}\rightarrow\nu_{2}$, and $\nu_{3}^{m}\rightarrow\nu_{3}$). In the scenario in which at least two of the matter eigenstates are produced in similar numbers, as described in the last paragraph, the possible range of flavor ratios at the Earth can be constrained to be $f_{\nu_{e}}^{NO}\lesssim 0.5$ and $f_{\nu_{e}}^{IO}\approx 1/3$. This result corresponds to the red-shaded region in Fig. 1. Notably, the IO scenario indicates a near equipartition among flavors leaving the SN, irrespective of the flavor output from self-induced conversions.
To better understand the emergence of these constraints, we characterize the evolution within the zone of standard matter effects in terms of flavor ratios $(f_{\nu_{e}},f_{\nu_{\mu}},f_{\nu_{\tau}})$. Assuming NO and adiabatic propagation, it follows that an initially produced $\nu_{e}$ arrives at the Earth as $\nu_{3}$ Dev:2023znd :
$(1,0,0)_{SN}\rightarrow\left(\left|U_{e3}\right|^{2},\left|U_{\mu 3}\right|^{2},\left|U_{\tau 3}\right|^{2}\right)_{\oplus}~{}.$ (7)
Conversely, either of the non-electron flavors $\nu_{\mu}$ or $\nu_{\tau}$ arrive at the Earth approximately as equal mixtures of $\nu_{1}$ and $\nu_{2}$. Therefore,
$\hskip 14.22636pt\hskip 14.22636pt\hskip 14.22636pt(0,1,0)_{SN}\hskip 14.22636pt\text{or}\hskip 14.22636pt(0,0,1)_{SN}\rightarrow$
$\frac{1}{2}\left(\left|U_{e1}\right|^{2}+\left|U_{e2}\right|^{2},\left|U_{\mu 1}\right|^{2}+\left|U_{\mu 2}\right|^{2},\left|U_{\tau 1}\right|^{2}+\left|U_{\tau 2}\right|^{2}\right)_{\oplus}~{}.$ (8)
For any initial combination $(a,b,c)_{SN}$, we obtain on Earth the $\nu_{e}$ fraction
$f_{\nu_{e}}^{NO}=a\left|U_{e3}\right|^{2}+b\frac{\left|U_{e1}\right|^{2}+\left|U_{e2}\right|^{2}}{2}+c\frac{\left|U_{e1}\right|^{2}+\left|U_{e2}\right|^{2}}{2}~{}.$ (9)
Because of the unitarity of the PMNS matrix and $a+b+c=1$, Eq. 9 simplifies to
$f_{\nu_{e}}^{NO}=\frac{1}{2}\left(1-\left|U_{e3}\right|^{2}\right)+\frac{a}{2}\left(3\left|U_{e3}\right|^{2}-1\right)~{},$ (10)
Adopting $\left|U_{e3}\right|^{2}\approx 0.02\ll 1$ deSalas:2020pgw , we obtain
$f_{\nu_{e}}^{NO}\approx\frac{(1-a)}{2}\lesssim 0.5~{}.$ (11)
In the IO scenario, the argument mirrors the one between Eq. 7 and Eq. 10, with the interchange of the PMNS matrix elements $U_{\alpha 3}$ and $U_{\alpha 2}$ for any flavor $\alpha$. Thus, by replacing $U_{e3}$ with $U_{e2}$ in Eq. 10, we have
$f_{\nu_{e}}^{IO}=\frac{1}{2}\left(1-\left|U_{e2}\right|^{2}\right)+\frac{a}{2}\left(3\left|U_{e2}\right|^{2}-1\right)~{}.$ (12)
Assuming $\left|U_{e2}\right|^{2}\approx 1/3$ deSalas:2020pgw , we obtain [^4] Eqs. 12 and 13 also have some curious implications in the context of solar neutrinos. Namely, they imply that any initial neutrino flavor configuration in the solar core leads to $f_{\nu_{e}}\approx 1/3$ at the Earth. This fact applies to both orderings, as $\nu_{e}$ is produced as $\nu_{2}^{m}$ and follows adiabatic propagation in the Sun’s interior.
$f_{\nu_{e}}^{IO}\approx\frac{1}{3}~{}.$ (13)
The red-shaded regions in Fig. 1 correspond to $f_{\nu_{e}}^{NO}$ and $f_{\nu_{e}}^{IO}$ after varying $\theta_{12}$ and $\theta_{13}$ within their $3\sigma$ allowed range and using the best-fit values for the other oscillation parameters (we have verified that the impact of $\theta_{23}$ and $\delta_{CP}$ on the resulting $f_{\nu_{e}}$ range is less than $10\%$) deSalas:2020pgw . The uncertainty in $f_{\nu_{e}}^{IO}$ arises predominantly from the uncertainty in the oscillation parameters, rather than the initial composition $(a,b,c)_{SN}$. As the precision in the determination of the oscillation parameters increases, these constraints should improve.
The results in Eq. 11 and Eq. 13 are applicable to neutrinos of any energy and at any instant in time (for the results related to antineutrinos, see the Appendix). Nonetheless, exceptions may be possible when a shock wave is present, leading to violations of Eq. 11 at certain parts of the energy spectrum.
Results in the presence of shock waves.— When adiabatic propagation cannot be guaranteed, transitions between different matter eigenstates may occur at the $H$- and $L$-resonance layers. Furthermore, if the matter profile is affected by shock waves, multiple resonances of the $H$- and $L$-types could appear Schirato:2002tg ; Tomas:2004gr . If an even number of $H$-resonances are non-adiabatic, results do not change [^5] Here we are neglecting phase effects that develop between a pair of shock waves, as the observation of these effects should be challenging in realistic detectors Dasgupta:2005wn .. However, if an odd number of the $H$-resonances are non-adiabatic, the outcome is modified in the NO scenario once transitions $\nu_{3}^{m}\leftrightarrow\nu_{2}^{m}$ that happen at one of the resonance layers are not reverted by another transition of the same type. It is straightforward to determine the consequence of an odd number of non-adiabatic $H$-resonances: transitions of the type $\nu_{3}^{m}\leftrightarrow\nu_{2}^{m}$ imply that matrix elements $U_{\alpha 3}$ and $U_{\alpha 2}$ should be interchanged in Eqs. 7-10. The argument is similar to the one that leads to Eqs. 12 and Eq. 13, resulting in
$f_{\nu_{e}}^{NO}(\text{non-ad $H$})\approx\frac{1}{3}~{}.$ (14)
Because the position of the $H$-resonance is energy dependent, the portion of the energy spectrum influenced by the transient non-adiabaticity provides information about the location of the shock wave front Schirato:2002tg ; Tomas:2004gr . Thus, we can track the motion of the shock wave through the $H$-resonance layers using flavor ratios if neutrino masses follow NO: during the shock’s transit through that region, $f_{\nu_{e}}$ converges to $1/3$ in the corresponding part of the energy spectrum.
In the IO scenario, the $H$-resonance does not occur in the neutrino channel. Therefore, the result is identical to the adiabatic case:
$f_{\nu_{e}}^{IO}(\text{non-ad $H$})\approx\frac{1}{3}~{}.$ (15)
[…]
Although shock wave fronts can also penetrate the $L$-resonance layer, adiabaticity is never strongly violated there Mirizzi:2015eza . Even in cases where such violations may happen, they typically occur at later times (beyond $10$ seconds post-bounce) Tomas:2004gr , so we do not consider them here.
Wave-packet decoherence during self-induced conversions.— All the derived results so far relied on the assumption that the off-diagonal elements in Eq. 6 are negligible. However, there are physical situations of interest in which Eq. 6 is not diagonal. For instance, the outcome of collective effects could be eigenstates in neutrino-neutrino potentials that adiabatically transform to matter eigenstates in the usual matter potentials. In such a situation, the density matrix for $N$ neutrinos would be diagonal in the matter eigenstates basis, but not in the flavor basis Akhmedov:2017mcc . The matter eigenstates $\nu_{1}^{m}$, $\nu_{2}^{m}$ and $\nu_{3}^{m}$ would arise in any relative amounts with no correlations between any two of them, and results would differ drastically from previous sections.
Another interesting situation would happen if some unknown physical mechanism restricts the allowed values of all parameters in Eq. 6. In this case, the fractions of $\nu_{1}^{m}$, $\nu_{2}^{m}$ and $\nu_{3}^{m}$ that enter the region of standard matter effects are strongly dependent on the specific values of $\alpha_{k}$, $\beta_{k}$ and $\phi_{k}$. If the range of values for these parameters is unknown or varies significantly over time, the initial relative fractions of each matter eigenstate would also appear to be arbitrary.
Within the scenarios described in this section, the outcome of collective effects is better formulated in terms of matter eigenstates rather than flavor eigenstates, and it resembles the usual picture of wave-packet decoherence of neutrino oscillations in vacuum and matter Porto-Silva:2021ael . Nevertheless, we can still constrain the flavor composition on Earth in terms of $f_{\nu_{e}}$. By considering these eigenstates to leave the collective effects region and, subsequently, arrive at Earth in arbitrary relative proportions, we find that $f_{\nu_{e}}\lesssim 0.7$ for both mass orderings [^6] Note that this constraint extends to hydrodynamic turbulence affecting neutrino propagation Mukhopadhyay:2023tsc , as well as other alternative scenarios that randomize the quantities of each of the vacuum eigenstates ($\nu_{1}$, $\nu_{2}$ and $\nu_{3}$) reaching the Earth.. These constraints are depicted as pink-shaded regions in the flavor triangles of Fig. 1.
Discussion and Conclusion.— We have shown that the flavor composition of SN neutrinos reaching the Earth can be constrained, even when assuming complete uncertainty about the neutrino state that emerges from the innermost regions dominated by self-induced conversions. Under certain conditions, the neutrino flux can be viewed as composed of flavor eigenstates in an arbitrary flavor ratio configuration (see Eq. 6 and related discussion). If this scenario occurs, we predict that the proportion of $\nu_{e}$ is always smaller than $0.5$ for NO, whereas, for IO, we expect near flavor equipartition ($f_{\nu_{e}}\approx 1/3$). These results are valid for neutrinos throughout the energy spectrum and duration of the emission (for the results related to antineutrinos, see the Appendix). If the $H$-resonance is non-adiabatic, for example, due to shock waves, flavor equipartition is the expectation for both mass orderings. However, in the opposite case in which neutrinos come out from self-induced conversions as fluxes of matter eigenstates (for instance, due to wave-packet decoherence), we obtain $f_{\nu_{e}}\lesssim 0.7$.
Ultimately, this Letter advocates for a phenomenological approach to comprehensively address the intricacies of neutrino flavor conversion in SN environments. Experimental challenges along the way include distinguishing between $\nu_{x}$ and $\bar{\nu}_{x}$ in neutral current scatterings Jachowicz:2004we . This approach can also be extended to accommodate non-standard scenarios, potentially unveiling novel phenomena underlying the physics of neutrino masses and mixing Jana:2022tsa ; Jana:2023ufy .
Acknowledgments.— We thank Orlando Peres for useful discussions. AC thanks for the support received by the scholarships CAPES/PROEX No. 88887.511843/2020-00, CNPq No. 140316/2021-3, CAPES-PrInt No. 88887.717489/2022-00 and FAPERJ No. E-26/204.138/2022. The work of YP was supported by the São Paulo Research Foundation (FAPESP) Grant No. 2023/10734-3 and 2023/01467-1, and by the National Council for Scientific and Technological Development (CNPq) Grant No. 151168/2023-7. MMS thanks for the support provided by iTHEMS and ABBL from RIKEN, the N3AS Physics Frontier Center, and the NSF under cooperative agreement 2020275. This work would not have been possible without the collaborative environment provided by the Nordic Winter School on Multimessenger Astrophysics. We thank the event organizers, as well as the financial support by Markus Ahlers for AC to attend the event.
Appendix
Antineutrinos.— For antineutrinos, the evolution Hamiltonian is similar to Eq. 1, with the modification $V_{e}\rightarrow-V_{e}$. Therefore, at $\rho_{m}\gg\rho_{m}^{H}$, $\bar{\nu}_{e}\approx\bar{\nu}_{1}^{m}$ for NO, and the possible fraction of electron antineutrinos on Earth is given by Eq. 10 with $\left|U_{e1}\right|^{2}$ in place of $\left|U_{e3}\right|^{2}$,
$f_{\bar{\nu}_{e}}^{NO}=\frac{1}{2}\left(1-\left|U_{e1}\right|^{2}\right)+\frac{a}{2}\left(3\left|U_{e1}\right|^{2}-1\right)~{}.$ (16)
Assuming $\left|U_{e1}\right|^{2}\approx 2/3$ deSalas:2020pgw , we have
$\frac{1}{6}\lesssim f_{\bar{\nu}_{e}}^{NO}\lesssim\frac{2}{3}~{}.$ (17)
Eq. 17 is applicable even in the presence of shock waves. On the other hand, for IO, $\bar{\nu}_{e}\approx\bar{\nu}_{3}^{m}$ above the $H$-resonance density, and $f_{\bar{\nu}_{e}}^{IO}$ is described by Eq. 10,
$f_{\bar{\nu}_{e}}^{IO}\lesssim 0.5~{}.$ (18)
If the $H$-resonance is non-adiabatic, transitions $\bar{\nu}_{3}^{m}\rightarrow\bar{\nu}_{1}^{m}$ modify the allowed regions for $f_{\bar{\nu}_{e}}^{IO}$, which will be given by Eq. 16. Therefore,
$\frac{1}{6}\lesssim f_{\bar{\nu}_{e}}^{IO}(\text{non-ad $H$})\lesssim\frac{2}{3}~{}.$ (19)
In case antineutrinos emerge from the core as incoherent matter eigenstates, we have $f_{\bar{\nu}_{e}}\lesssim 0.7$ for both mass orderings. These results are depicted in Fig. 2.
Figure 2: Same as Fig. 1, but for antineutrinos.